A parachutist's rate during a free fall reaches 231 feet per second. What is this rate in meters per second? At this rate, how many meters will the parachutist fall during 15 seconds of free fall? In your computations, assume that 1 meter is equal to 3.3 feet. Do not round your answers.
1
Expert's answer
2020-02-24T11:08:12-0500

"1\:meter\:=\:3.3\:feets"

"\Rightarrow" "1\:meter\:per\:second\:=\:3.3\:feet\:per\:second"


Rate in meter per second will be given by

"231\:feet\:per\:second\:=\:231\:feet\:per\:second\:\cdot \frac{1\:meter\:per\:second}{3.3\:feet\:per\:second}"

"=\frac{231}{3.3}\:meters\:per\:second"

"=70 \:meters\:per\:second"


Distance = Rate * time

In 15 seconds, parachutist will fall = "70\:\frac{meter}{second\:}\cdot 15\:seconds= 1050\: meters"